From nch@roe.ac.uk Mon Aug 30 11:39:22 2010 Date: Fri, 27 Aug 2010 16:05:03 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren , Tom Kerr Subject: WFAU VDFS Science Archives fortnightly meet mins 27/08/10 Minutes of WFAU VDFS Science Archive meeting: August 27th 2010 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NJC, MAR, NCH, RSC, ETWS, RPB, MSH Apologies: JPE, AL, RGM, KTN DoNM: Friday September 10th in the Vista Hut Actions discharged this week: ----------------------------- ACTION: RSC & NJC to document new operational procedures on the TWiki Discharged; now up to date with procedures as at DR8. ACTIONS: - Change schema / CU4 for new aperture mag names and columns for corrections (partly discharged by ETWS and to be finalised by NJC) - Updating Filter table with A/E(B-V) and VegaToAB (Need values from STH) Continues Discharged by NJC and ETWS. ACTION: RSC to add new ppErrBit to GPS catalogues as per instructions from the survey head; and deprecate affected framesets to ensure correct propagation/reseaming in gpsSource DR7. Discharged Actions partly discharged but continuing: ----------------------------------------- ACTION: NCH to assign actions on the following next time: - Modify entire code base to work with Tiles (Partly discharged - but needs fully testing on the new V1 data). This includes checking code works here on deep stacks, checking quality control/ QBF, ingest, CU7, CU6. Also do we need to do better links between tile detections and pawprint detections or will Provenance etc be good enough. (RSC has added tile qbits; also suggested a preliminary test run of CU7 using shallow tiles first of all). - VVV SynopticSource (allowing both SourceXSynopticSourceBestMatch and SourceXDetectionBestMarch) ACTION: MAR to add some web documentation for the GPS photometric calibration issues. Actions carried forward from 13/08/10 meeting --------------------------------------------- ACTION: TBD to draft and MOU on behalf of PSPIs for ESO-SAF deliveries Continues; on hold until at least next PSPI teleconf, and possibly until "Phase 3" workshop at ESO. ACTION: NJC to suggest workshop at next PSPI meeting. NJC noted that he needs to liaise with JPE about the next teleconf. ACTION: RPB to update WFCAMOPENTIME for 09A projects no longer inside their proprietary period. RPB noted that he will check this. Specific points and new actions: -------------------------------- Project management: Nothing new this week. WFCAM & VISTA updates: No updates this week. Comments and issues arising from CASU minutes: No new minutes this week. Networking: WFCAM data transfers are all up to date for 10A; latest batch of VIRCam data (up to July 23rd) has become available from CASU today so transfers will continue. WSA/VSA Operations: ETWS noted -- Filled up the newly added attributes with all available data from the files. -- Finished transfer and metadata ingest for WFCAM data up to July 31st. -- VISTA catalogue data ingest had some holdup due to disk space problems and dodgy data. I've included a new check that should take care of the latter. The disk space problem was due to producing ingestable catalogues quicker (since it is parallelised) than they could be ingested. -- Processed WFCAM surveys up to the final release curation task. DR8 preparations are all complete (following a few hiccups introduced by changes required for VSA curation) and the release DB is in the process of being published. MAR and NJC noted that a VSA QC issue has come up regarding unplanned repeats in tile areas due to aborted/incomplete OBs. It's likely that a new deprecation code is need to flag (at least) those members of aborted MSBs. Hardware and Systems: It was agreed that a new cluster node should be acquired with which to host a separate VVV ingest DB since observations during the middle of the year are totally dominated by VVV catalogues, and it's holding up progress for VISTA surveys. MSH noted that we know the required spec and it's simply a matter of placing on order. Software: RSC noted: "A new detection quality flag for poorly calibrated WFCAM GPS frames in the Galactic centre region has been implemented and applied to latest merged source table for DR7. It is documented on the website and in the forthcoming data releases' schema glossary. We await news from Phil Lucas as to what criterion we should use to apply a similar flag to the VISTA VVV data, but will release the data in the meantime if it is ready before the flag can be implemented. Steve Warren has noticed that the quality flag for detections in the underexposed dither region of WFCAM LAS stacks was slightly too conservative in that not every detection with poor photometry has been flagged. We've asked for advice on how best to modify the existing implementation to ensure these detections are flagged for DR9." Survey Data Release: UKIDSS DR8 is busily copying out from the WSA ingest DB for deployment later today on the public DB servers. UKIDSS Survey Head SJW scared the cr*p out of everyone earlier in the week claiming that he'd spotted a "possible serious calibration issue" but the effect in question turned out to be astrophysical in origin - ye canae change the laws of physics, cap'n! It was agreed to pause general VSA ingests to build release DBs for the VHS and VMC including data up to the end of March. Non-survey Data Release: Nothing of note this week. Astrogrid deployment & Data Analysis services: MSH noted that he is upgrading the WFAU DSAs with ADQL2 which will (amongst other things) enable federation with VO datasets over Table Access Protocol. New functionality includes generalised crossmatching facilities of arbitrary VO-published catalogues. Miscellaneous: That's all folks. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================